rewrite 2x-3y=-6 as a function of f (x)
Answer by stanbon(75887)

rewrite 2x-3y=-6 as a function of f (x)
3y = 2x+6
y = (2/3)x + 2
===================
If you want it done as a function of y you get:
2x = 3y-6
x = (3/2)y-3
=================
Cheers,
Stan H.